# Sequence-Augmented $\mathrm{SE(3)}$-Flow Matching For Conditional Protein Backbone Generation

### v1

@@ Abstract
Proteins are essential for almost all biological processes and derive their diverse functions from complex $3\rm D$ structures, which are in turn determined by their amino acid sequences. In this paper, we exploit the rich biological inductive bias of amino acid sequences and introduce FoldFlow-2, a novel sequence-conditioned $\mathrm{SE(3)}$-equivariant flow matching model for protein structure generation. FoldFlow-2 presents substantial new architectural features over the previous FoldFlow family of models including a protein large language model to encode sequence, a new multi-modal fusion trunk that combines structure and sequence representations, and a geometric transformer based decoder. To increase diversity and novelty of generated samples—crucial for de-novo drug design—we train FoldFlow-2 at scale on a new dataset that is an order of magnitude larger than PDB datasets of prior works, containing both known proteins in PDB and high-quality synthetic structures achieved through filtering. We further demonstrate the ability to align FoldFlow-2 to arbitrary rewards, e.g. increasing secondary structures diversity, by introducing a Reinforced Finetuning (ReFT) objective. We empirically observe that FoldFlow-2 outperforms previous state-of-the-art protein structure-based generative models, improving over RFDiffusion in terms of unconditional generation across all metrics including designability, diversity, and novelty across all protein lengths, as well as exhibiting generalization on the task of equilibrium conformation sampling. Finally, we demonstrate that a fine-tuned FoldFlow-2 makes progress on challenging conditional design tasks such as designing scaffolds for the VHH nanobody.
@@ 3.2 Training
During training, we set the fraction of synthetic samples that may be seen during an epoch to $2/3$ of the epoch. This prevents the model from overfitting to the remaining noise in the synthetic data, and is also common practice when training with synthetic data (Hsu et al., 2022; Lin et al., 2023). Anecdotally, we did not notice an improvement from using a smaller proportion of synthetic structures. Finally, in the FoldFlow-2 architecture, we keep the ESM pre-trained language model fixed during training and train all other components (encoder, trunk, and decoder) from scratch. The results presented in 3 use PDB data only, as this displayed the best performance for designability scores.
@@ 4.4 Motif Scaffolding
Model Solved /24 Motif Scaffold Solved /25
RFDiffusion 24 3.94 $\pm$ 1.54 2.40 $\pm$ 0.93 5
@@ 4.5 Zero-shot Equilibrium Conformation Sampling
We now test FoldFlow-2 on zero-shot equilibrium conformation sampling task. Starting from a sequence, we generate multiple conformations of the same proteins and compare the distribution of conformations with the ones from molecular dynamic simulations. We compare FoldFlow-2 with AlphaFlow-MD and ESMFlow-MD; two folding models fine-tuned on a molecular dynamic dataset. In table 6, we report the pairwise and global RMSD, the root mean square fluctuation (RMSF), and the 2-Wasserstein on the top two principal components. For both RMSD and the RMSF metrics, we report the Pearson correlation between the values from the generated ensemble and those of the ground truth ensemble (the procedure is detailed in full in §B.7).
@@
Table 6: Zero-shot performance of the base FoldFlow-2 model on the ATLAS dataset of MD trajectories compared to ESMFlow and AlphaFlow models fine-tuned on ATLAS. FoldFlow-2 is competitive to the comparable model ESMFlow across all metrics. $r$ denotes Pearson’s correlation coefficient.
Model Pairwise RMSD $r$ ($\uparrow$) Global RMSF $r$ ($\uparrow$) Per-target RMSF $r$ ($\uparrow$) PCA $\mathcal{W}_{2}$-dist ($\downarrow$)
AlphaFlow-MD 0.468 $\pm$ 0.005 0.415 $\pm$ 0.006 0.824 $\pm$ 0.000 10.67 $\pm$ 0.29
ESMFlow-MD 0.293 $\pm$ 0.005 0.161 $\pm$ 0.005 0.737 $\pm$ 0.001 11.51 $\pm$ 0.13
FoldFlow-2 0.297 $\pm$ 0.004 0.236 $\pm$ 0.004 0.658 $\pm$ 0.001 10.85 $\pm$ 0.15
@@ 5 Related work
Protein design. Physics-based protein structure design yielded the first de novo proteins (Huang et al., 2016). For example, structure-based biophysics approaches have previously resulted in several drug candidates (Röthlisberger et al., 2008; Fleishman et al., 2011; Cao et al., 2020).This was followed by language models (Hie et al., 2022; Ferruz et al., 2022) and geometric deep learning (Gainza et al., 2020) for protein structure design. Recently, diffusion (Wu et al., 2024; Yim et al., 2023b; Watson et al., 2023; Ingraham et al., 2023; Wang et al., 2024; Frey et al., 2023) and flow-based models (Bose et al., 2024; Yim et al., 2023a; Jing et al., 2024) have risen to prominence. These methods employ a backbone-first approach with the notable exception of MultiFlow (Campbell et al., 2024) which uses sequence to perform co-generation.
@@ B.6 Motif scaffolding
1BCF 98 98 98 98 100
2KL8 95 95 95 95 100
1PRW 77 99 99 77 91
6EXZ_long 70 80 82 68 91
6EXZ_med 60 73 76 60 87
1QJG 73 62 93 54 80
5TRV_long 48 49 85 36 77
5TRV_med 50 46 84 33 80
6E6R_long 35 72 95 31 82
6EXZ_small 30 41 49 30 28
6E6R_med 28 84 95 27 87
4ZYP 36 53 87 25 85
6E6R_small 24 81 97 21 50
1YCR 22 91 98 20 91
7MRX_small 18 31 67 15 22
5TPN 14 32 59 14 79
5TRV_small 31 22 75 13 53
3IXT 14 89 93 13 85
5YUI 24 12 40 9 8
7MRX_med 17 32 66 6 22
7MRX_long 15 21 68 4 22
5IUS 15 4 56 3 11
4JHW 12 3 38 1 8
5WN9 1 6 9 1 1

### v2

@@ Abstract
Proteins are essential for almost all biological processes and derive their diverse functions from complex $3\rm D$ structures, which are in turn determined by their amino acid sequences. In this paper, we exploit the rich biological inductive bias of amino acid sequences and introduce FoldFlow-2  [^1] Our code can be found at https://github.com/DreamFold/FoldFlow, a novel sequence-conditioned $\mathrm{SE(3)}$-equivariant flow matching model for protein structure generation. FoldFlow-2 presents substantial new architectural features over the previous FoldFlow family of models including a protein large language model to encode sequence, a new multi-modal fusion trunk that combines structure and sequence representations, and a geometric transformer based decoder. To increase diversity and novelty of generated samples—crucial for de-novo drug design—we train FoldFlow-2 at scale on a new dataset that is an order of magnitude larger than PDB datasets of prior works, containing both known proteins in PDB and high-quality synthetic structures achieved through filtering. We further demonstrate the ability to align FoldFlow-2 to arbitrary rewards, e.g. increasing secondary structures diversity, by introducing a Reinforced Finetuning (ReFT) objective. We empirically observe that FoldFlow-2 outperforms previous state-of-the-art protein structure-based generative models, improving over RFDiffusion in terms of unconditional generation across all metrics including designability, diversity, and novelty across all protein lengths, as well as exhibiting generalization on the task of equilibrium conformation sampling. Finally, we demonstrate that a fine-tuned FoldFlow-2 makes progress on challenging conditional design tasks such as designing scaffolds for the VHH nanobody.
@@ 3.2 Training
During training, we set the fraction of synthetic samples that may be seen during an epoch to $2/3$ of the epoch. This prevents the model from overfitting to the remaining noise in the synthetic data, and is also common practice when training with synthetic data (Hsu et al., 2022; Lin et al., 2023). Anecdotally, we did not notice an improvement from using a smaller proportion of synthetic structures. Finally, in the FoldFlow-2 architecture, we keep the ESM pre-trained language model fixed during training and train all other components (encoder, trunk, and decoder) from scratch. The results presented in table 3 and §4.4 use PDB data only, as this displayed the best performance for designability scores.
@@ 4.4 Motif Scaffolding
Model Solved /24 $\uparrow$ Diversity $\uparrow$ Motif $\downarrow$ Scaffold $\downarrow$ Solved /25 $\uparrow$
RFDiffusion 24 0.345 3.94 $\pm$ 1.54 2.40 $\pm$ 0.93 5
@@ 4.5 Zero-shot Equilibrium Conformation Sampling
We now test FoldFlow-2 on zero-shot equilibrium conformation sampling task. Starting from a sequence, we generate multiple conformations of the same proteins and compare the distribution of conformations with the ones from molecular dynamic simulations. We compare FoldFlow-2 with AlphaFlow-MD and ESMFlow-MD; two folding models fine-tuned on a molecular dynamic dataset, and non finetuned models Eigenfold (Jing et al., ) and STR2STR (Lu et al., 2024). In table 6, we report the pairwise and global RMSD, the root mean square fluctuation (RMSF), and the 2-Wasserstein on the top two principal components. For both RMSD and the RMSF metrics, we report the Pearson correlation between the values from the generated ensemble and those of the ground truth ensemble (the procedure is detailed in full in §B.7).
@@
[…]
AlphaFlow-MD 0.468 $\pm$ 0.005 0.415 $\pm$ 0.006 0.824 $\pm$ 0.000 10.67 $\pm$ 0.29 32.6 $\pm$ 0.1
ESMFlow-MD 0.293 $\pm$ 0.005 0.161 $\pm$ 0.005 0.737 $\pm$ 0.001 11.51 $\pm$ 0.13 11.2 $\pm$ 0.1
FoldFlow-2 0.297 $\pm$ 0.004 0.236 $\pm$ 0.004 0.658 $\pm$ 0.001 10.85 $\pm$ 0.15 5.273 $\pm$ 0.075 9.0 $\pm$ 0.0
STR2STR 0.279 $\pm$ 0.003 0.244 $\pm$ 0.005 0.586 $\pm$ 0.001 5.37 $\pm$ 0.05 2.678 $\pm$ 0.023 10.6 $\pm$
Eigenfold 0.194 $\pm$ 0.004 0.156 $\pm$ 0.004 0.668 $\pm$ 0.001 5.296 $\pm$ 0.080 38.6 $\pm$ 0.5
@@ 5 Related work
Protein design. Physics-based protein structure design yielded the first de novo proteins (Huang et al., 2016). For example, structure-based biophysics approaches have previously resulted in several drug candidates (Röthlisberger et al., 2008; Fleishman et al., 2011; Cao et al., 2020).This was followed by language models (Hie et al., 2022; Ferruz et al., 2022) and geometric deep learning (Gainza et al., 2020) for protein structure design. Recently, diffusion (Wu et al., 2024; Yim et al., 2023b; Watson et al., 2023; Ingraham et al., 2023; Wang et al., 2024; Frey et al., 2024) and flow-based models (Bose et al., 2024; Yim et al., 2023a; Jing et al., 2024) have risen to prominence. These methods employ a backbone-first approach with the notable exception of MultiFlow (Campbell et al., 2024) which uses sequence to perform co-generation.
@@ B.6 Motif scaffolding
1BCF 100 100 100 100 100
2KL8 100 100 100 100 100
1PRW 100 100 98 98 91
1YCR 96 100 89 85 91
4ZYP 89 97 87 80 85
3IXT 97 101 75 73 85
7MRX_small 70 92 83 66 22
6EXZ_long 78 83 62 59 91
5TPN 62 95 78 57 79
6EXZ_med 67 70 55 54 87
6EXZ_small 56 63 55 53 28
6E6R_long 84 98 57 51 82
1QJG 54 82 76 49 80
7MRX_med 59 86 67 49 22
6E6R_med 78 96 54 43 87
5TRV_med 44 86 70 41 80
5TRV_long 40 85 69 38 77
6E6R_small 82 100 34 30 50
5TRV_small 28 78 47 24 53
7MRX_long 27 59 47 20 22
5YUI 14 54 48 13 8
5IUS 10 86 79 10 11
5WN9 6 12 10 4 1
4JHW 3 74 39 3 8